OpenStudy (anonymous):

Simplify \(\frac{cos^{2}a}{1-tana} + \frac{sin^{2}a}{1-cota}\)

OpenStudy (helder_edwin):

\[\large \frac{\cos^2a}{1-\tan a}+\frac{\sin^2a}{1-\cot a}= \frac{\cos^2a}{1-\tan a}+\frac{\sin^2a}{1-\frac{1}{\tan a}}= \] \[\large =\frac{\cos^2a}{1-\tan a}+\frac{(\tan a)(\sin^2a)}{\tan a-1}= \frac{\cos^2a}{1-\tan a}-\frac{(\tan a)(\sin^2 a)}{1-\tan a} \]

OpenStudy (anonymous):

but you should Simplify it

OpenStudy (helder_edwin):

\[\large =\frac{\cos^2a-(\tan a)(\sin^2a)}{1-\tan a}= \frac{\cos^2a-(\tan a)(1-\cos^2a)}{1-\tan a}= \] \[\large =\frac{\cos^2a-\tan a+(\tan a)(\cos^2a)}{1-\tan a} \]

OpenStudy (helder_edwin):

i know.that is what i was trying to do